1 a flywheel rotating about fix axis  has a kinetic energy of 360 joule when it angular speed 30 Radian per second the moment of inertia of wheel

Grade:12th pass

2 Answers

Vikas TU
4 years ago
Hiii 
Moment of Inertia = Mr^2
omega = 30 radian , and kinetic energy = 360 joule
Rotational KE = ½ * I * W^2 ( W = omega)
Put the value you will get 
I = 12/15 = 0.80
Khimraj
4 years ago
I = mrr
moment of inertia is equal to the product of mass and the square of the radius.
we know rotational kinetic energy = 1/2iww.
then given w= 30 rad and k.E = 360j 
k.e= 1/2iww
 360 = 1/2× i×30×30
i= 360×2÷30×30
I=12÷15